OpenStudy (anonymous):

A fair coin is tossed twice. what is the probability of obtaining at least one head?

OpenStudy (dumbcow):

There are 4 possibilities HH HT TH TT 3 of these have at least one H ->P = 3/4

OpenStudy (anonymous):

You could also recognize, P(at least 1 head) = 1 - P(no heads) = 1 - P( 2 tails) = 1 - 1/2 x 1/2 = 1-1/4 = 3/4

OpenStudy (anonymous):

What lokisan is doing is finding the probability that you will get tails twice, (which would be multiplying one half by one half), then subtracting that from one, which would get the probability that you wouldn't get tails twice, which is the same thing as getting heads at least once.